bella2025:

2. Solve the system by elimination. Hint: you will get a decimal for one of your coordinates. 4x - 5y = -9 -4x + y = -7 -5y=-9 1y=-7 -4y= -16 y=4 -4x+4=-7 -4x=-11 x=2.75 @axie or anyone did I do this correctly

axie:

-5y=-9 1y=-7 can you explain what you did to convert this? i.e. How you got from -5y to 1y

bella2025:

@axie wrote:
-5y=-9 1y=-7 can you explain what you did to convert this? i.e. How you got from -5y to 1y
4x - 5y = -9 -4x + y = -7 so the x's cancel out because you have 4x and -4x so then your left with -5y=-9 and -y=-7 because if you have a variable by itslef you put a 1 infront of it

axie:

Wouldn't the one be positive bc the negative cancelled out when you combined the x's. So I think it would be -5y=-9 AND 1y=-7 then you would combine those two equations. Again i might be wrong @timmyspu is that correct?

Timmyspu:

You are doing it correct @axie

bella2025:

@axie wrote:
Wouldn't the one be positive bc the negative cancelled out when you combined the x's. So I think it would be -5y=-9 AND 1y=-7 then you would combine those two equations. Again i might be wrong @timmyspu is that correct?
and I did combine them-5 plus 1 is -4y and -9+-7=-16

axie:

OH ok, then you devided both sides by 4 which gave you y=4

bella2025:

@axie wrote:
OH ok, then you devided both sides by 4 which gave you y=4
yeah and the I plugged 4 into y thats how I got -4x+4=-7

axie:

Ok, so both equations are correct, I was a bit lost, then again i do my math differently, good job!

Timmyspu:

Way to go @axie and @bella2025